OpenStudy (anonymous):

(2a+5) cube of a binomial

OpenStudy (idku):

\[(a+b)^3=(a+b)(a^2+ab+b^2)=a^3+2a^2b+2ab^2+b^3\]

OpenStudy (idku):

your "a" is 2a, and your "b" is 5 .